A random sample of vehicle mileage expectancies has a sample mean of x¯=169,200 miles and sample standard deviation of s=19,400 miles. Use the Empirical Rule to estimate the percentage of vehicle mileage expectancies that are more than 188,600 miles.

Round your answer to the nearest whole number (percent).
Answer #1

Z = (X - xbar) / s = (188600 - 169200)/19400 = 1

P(X bar > 188600) = P(Z > 1) = (100 - 68) / 2 = 16%